At what angle shall a curve be raised in order for a 1500 kg vehicle to make a turn around a 100.0 m radius of a circular track at a velocity of 30.0 m/s?


1
Expert's answer
2021-11-25T10:10:21-0500

The angle of the banked curve is


"\theta=\arctan\frac{v^2}{rg}=42.6°."